# Retrieval of fundamental material parameters of monolayer transition metal dichalcogenides from experimental exciton energies: An analytical approach

### v1

@@ V Conclusion
This study proposes an analytical approach for retrieving fundamental material parameters such as the bandgap energy, screening length, exciton reduced mass, and surrounding dielectric constant from the experimental magnetoexciton energies in monolayer transition metal dichalcogenides (TMDCs). Unlike previous fitting methods, the approach presented here allows for the independent determination of the bandgap energy and screening length by the analytical formulas of the field-free exciton energies for $s$ states, with the fitting procedure is applied for only one parameter, the exciton reduced mass. The surrounding dielectric constant is then also determined through an analytical formula. By applying the suggested method to the currently available experimental data for exciton energies in monolayer TMDCs encapsulated by hBN, we retrieve fundamental material properties that complement and correct the existing data. The method is simple and easy to use for other experiments to retrieve fundamental material properties.
@@ Acknowledgments
V.-H.L., D.-N.Ly, and D.-N.Le conceptualized the work and performed analytical formulation; D.-N.Ly, D.-K.D.L., and V.-H.L. carried out numerical calculations; V.-H.L., D.-N.Ly, and D.-N.Le analyzed data, V.-H.L., D.-N.Ly, and D.-N.Le wrote and edited the manuscript.

### v2

@@ V Conclusion
We have successfully developed an analytical approach for retrieving fundamental material parameters, such as bandgap energy, screening length, exciton reduced mass, and surrounding dielectric constant, from experimental magnetoexciton energies in monolayer transition metal dichalcogenides (TMDCs). Unlike previous fitting methods that required three parameters, our approach allows for the independent determination of the bandgap energy and screening length from the field-free $s$-state exciton energies. This results in a one-parameter problem, with the exciton reduced mass being the only parameter that needs to be defined. Notably, for the first time, the bandgap energy and screening length can be computed directly from exciton energies using straightforward analytical formulas.
Furthermore, using perturbation theory combined with the Padé approximant, we have derived an analytical expression for magnetoexciton energies as functions of the exciton reduced mass. This allows for the extraction of the exciton reduced mass analytically from experimental data on magnetoexciton energies. The surrounding dielectric constant is then also determined through an analytical formula. Besides, we have also developed analytical expressions for diamagnetic coefficients and exciton radii using the extracted material data. This analytical theory for magnetoexciton energies is significant and valuable for analyzing experimental data.
By applying our method to the currently available experimental data for exciton energies in monolayer TMDCs encapsulated by hBN, we have retrieved fundamental material properties that complement and correct the existing data. The computation process is straightforward and user-friendly, making it suitable for other experiments aimed at retrieving fundamental material properties.
@@ Acknowledgments
Contribution: V.-H.L., D.-N.Ly, and D.-N.Le conceptualized the work and performed analytical formulation; D.-N.Ly, D.-K.D.L., and V.-H.L. carried out numerical calculations; V.-H.L., D.-N.Ly, D.-N.Le, and D.-K.D.L analyzed data, V.-H.L., D.-N.Ly, and D.-N.Le wrote and edited the manuscript.